Hawaii-2015-03 experiment with new MOBYrefresh blue Resonon - BS01 cfg002

First set of data after the BS01 came back from Resonon for some fixes.

Problems seen (and solved):

  1. Track Shoulder: There is still a shoulder on the side of tracks. Interestingly the shoulder is not seen or smaller when all the tracks are on. - seen on day 1 data
  2. 2nd Order Straylight: There is also the HUGE 2nd order bump on the 350 nm laser data (Page Num 19.01) removes the peaks, but is also removing too much of the UV light. So we are in search of a better filter.
  3. Another ghost image: The day 3 Es data showed a feature in the 0-30 pixel range which I have not seen before with the laser data. Mike Kehoe thinks it might be the infrared artifact. And suggested a laser at about 715-720 nm and 920 nm and 955 nm should generate big ghost images if the filter coating is the cause of the problem. When Mike put in laser data at those wavelengths I did not see the feature, See Page Num 12.02. Not sure what this means.
  4. Fiber modes: On day 4 we started testing the FO splitter. The day 4-6 data where not taken simultaneously on two tracks. On days 7-9 Mike collected simultaneous data on tracks 5 and 10 for Es head angles 0:10:90 to look for problems with fiber modes. There was a difference between the ratios, but when the data was collected again (day 9) keeping the light levels steady the problem went away. Which I think means the problem was stray light not fiber modes.
  5. Splitter Ratios: On day 10-12 we used laser data rather than lamp data to look at the FO splitter and found that the laser data ratios are not the same as the lamp ratios, when ratioing one pixel to the other with laser data. But when you integrate over +-5 pixeles the ratio do match up so problem solved. See Page Num 12.01 and 12.05.
  6. Lu Head Polarized: The day 13 data showed no polarization for the MOBY Es heads, but the day 14-15 data showed the Lu heads are polarized. When Mike added a depolarizer to the MOBY head this fixed the problem (See Page Num 17.01)
  7. Infrared Artifact: When the 720SP filter was added to Track 12 the infrared artifact was eliminated, Page Num 18.01
  8. Fluorescent light in the Laser: While taking laser data (Page Num 19.01) on days 19-20 with the LP and SP filters it was noticed that there was a large hump in the middle of the array for laser wavelengths 325 and 380 nm. Because of the filters this can not be strayligth so Ken best guess is fluorescent light from somewhere in the system (probably the laser?). Mike wonders if it might be sphere fluorescence.
  9. Vertical Pixel Shift Speed and Horizontal Pixel Shift Speed: See Page Num 20.01, 20.02 and Mikes PDF. Still making a decision.

LOG SHEETS

Photos

All the raw data files and their images , Table of KEYWORDS variable
Page Number
Link
Description
Date
1.01
Day 1 - 1 July Raw Data , Table
Laser data - 450,350 and 680 (all tracks, track 7 only, and tracks 4,7, and 10)
July 1, 2015
1.02
Day 2 - 2 July Raw Data , Table
Laser data - 450 nm - Track 7 only - looking at "Shutter open & close delay" time (msec)
July 2, 2015
1.03
Day 3 - 3 July Raw Data , Table
Es using a MOBY Ed head - track 7 only
July 2, 2015
1.04
Day 4 - 23 July - Raw Data, Table
Track 7 - FO splitter test - Oriel 60000 - fiber - problems - data repeated on Day 5
July 27, 2015
1.05
Day 5- 24 July - Raw Data, Table
Track 7 - FO splitter test - Oriel 60000 - fiber
July 27, 2015
1.06
Day 6- 29 July - Raw Data, Table
Track 7 - FO splitter test - Oriel 60000 - fiber and MOBY Es/Ed head
July 29, 2015
1.07
Day 7- 30 July - Raw Data, Table
Track 5/10 - FO splitter test -Oriel 60000 - MOBY Es/Ed head simultaneous w/ Rotation
July 29, 2015
1.08
Day 8- 1 Aug- Raw Data, Table
Track 5/10 - FO splitter test -Oriel 60000 - MOBY Es/Ed head simultaneous w/ Rotation
Aug 1, 2015
1.09
Day 9- 5 Aug- Raw Data, Table
Track 5/10 and 3/12- FO splitter test - Oriel 60000 - MOBY Es/Ed head simultaneou w/ Rotations - Repeat day 8 w/ dimmer lamp / higher int time
Aug 10, 2015
1.10
Day 10- 12 Aug- Raw Data, Table
Track 3/12- FO splitter test - Laser- MOBY Es/Ed head simultaneous w/ Rotation - 350, 375 and 400 nm
Aug 14, 2015
1.11
Day 11- 14 Aug- Raw Data, Table
Track 3/12- FO splitter test - Laser- MOBY Es/Ed head simultaneous w/ Rotation - 425, 450, 500 and 600 nm
Aug 14, 2015
1.12
Day 12- 14 Aug- Raw Data, Table
Track 3/12- FO splitter test - Laser- MOBY Es/Ed head simultaneous w/o Rotation - lots of laser wavelenth nm
Aug 17, 2015
1.13
Day 13- 14 Aug- Raw Data, Table
Track 3/12- FO splitter - Laser/Oriel Lamp- MOBY Es/Ed head simultaneous - Polarizer data
Aug 17, 2015
1.14
Day 14- 20 Aug- Raw Data, Table
Track 3/12- FO splitter - Oriel Lamp- MOBY Lu head simultaneous - Polarizer data = long FO
Aug 20, 2015
1.15
Day 15- 21 Aug- Raw Data, Table
Track 3/12- FO splitter - Oriel Lamp- MOBY Lu head simultaneous - Polarizer data = short FO
Aug 21, 2015
1.16
Day 16- 22 Aug- Raw Data, Table
Track 3/12- FO splitter - Oriel Lamp- MOBY Lu head simultaneous - 2 Polarizers crossed
Aug 22, 2015
1.17
Day 17- 25 Aug- Raw Data, Table
Track 3/12- FO splitter - Oriel Lamp- MOBY Lu head simultaneous - Polarizer data with Thorlabs Quartz Depolarizer behind Lu window
Aug 25, 2015
1.18
Day 18- 28 Aug- Raw Data, Table
Track 3/12- FO splitter - Oriel Lamp- MOBY Lu head - Carols long-pass & short-pass filters on Track 12
Aug 28, 2015
1.19
Day 19- 28 Aug- Raw Data, Table
Track 3/12- FO splitter - Laser - MOBY Lu head - Carols long-pass & short-pass filters on Track 12
Sept 1, 2015
1.20
Day 20- 28 Aug- Raw Data, Table
Track 3/12- FO splitter - Laser - MOBY Lu head - Carols long-pass & short-pass filters on Track 12
Sept 1, 2015
1.21
Day 21- 2 Sep- Raw Data, Table
Track 3/12- FO splitter - Laser - MOBY Lu head - LP and SP on Track 12 - different combinations of Horizontal readout rates and Vertical shift speeds
Sept 1, 2015
 
Day 1 - Laser data - 450,350 and 680 (all tracks, track 7 only, and tracks 4,7, and 10)
2.01
Track selection
Track selection using the day 1 laser data
July 1, 2015
2.02 Cross track shoulder Cross track shoulder seen in this data July 2, 2015
2.03 Wave Cal Wavelength calibration from the 1 day lasers July 6, 2015
2.04
Track selection - edge2edge
Track selection using the day 1 laser data - Edge to Edge tracks
Aug 19, 2015
Day 2 - Laser data - 450 nm - Track 7 only - looking at "Shutter open & close delay" time (msec)
3.01 CT and shutter Cross Track shoulder for track 7 verses shutter open & close delay July 2, 2015
Day 3 - Es using a MOBY Ed head - track 7 only
4.01 Es data A look at the Es data (cross and long track and shoulders) July 6, 2015
4.02 Es_data_vs_wave Same Es data but plotted verse wavelength - VERY ROUGH system response July 6, 2015
Day 4 - Track 7 - FO splitter test - Oriel 60000 - fiber - problems - data repeated on Day 5
5.01 FO Splitter FO Splitter Test - problem with weaker FO July 27, 2015
5.02 Cross Track Looking at Cross Track July 27, 2015
Day 5 - Track 7 - FO splitter test - Oriel 60000 - fiber
6.01 FO Splitter FO Splitter Test July 27, 2015
6.02 Ambient light leaks Looking for ambient light leaks (none found) July 27, 2015
6.03 Cross Track Looking at Cross Track July 28, 2015
6.04 Dark stability Looking at the Dark stability over day 4 and 5 data July 29, 2015
Day 6 - Track 7 - FO splitter test - Oriel 60000 - fiber and MOBY Es/Ed head
7.01 Fo Splitter FO Splitter Test - fiber data July 29, 2015
7.02 Fo splitter MOBY Es FO Splitter Test - MOBY Es head with rotation July 29, 2015
Day 7 - Track 5/10 - FO splitter test -Oriel 60000 - MOBY Es/Ed head simultaneous
8.01 Fo Splitter sim FO Splitter Test - fiber data - simultaneous splitter data July 30, 2015
8.02 Saddle shapes Looking for any changes in the saddle shapes July 31, 2015
Day 8 - Track 5/10- FO splitter test - Oriel 60000 - MOBY Es/Ed head simultaneous - repeat day 7
9.01 Fo Splitter sim FO Splitter Test - fiber data - simultaneous splitter data July 30, 2015
9.02 Saddle shapes Looking for any changes in the saddle shapes July 31, 2015
9.03 more saddle shapes Looking for any changes in the saddle shapes with more pixels Aug 3, 2015
Day 9 - Track 5/10 and 3/12 - FO splitter test - Oriel 60000 - MOBY Es/Ed head simultaneous - Repeat day 8 w/ dimmer lamp / higher int time
10.01 Fo Splitter sim Tracks 5/10 - fiber data - simultaneous splitter data Aug 11, 2015
10.02 PDF Mikes processing of the day 9 data Aug 11, 2015
10.03 Fo Splitter sim- higher int Tracks 5/10 - fiber data - simultaneous splitter data - int times increased to maintain light levels Aug 11, 2015
10.04 Fo Splitter sim Tracks 3/12 - fiber data - simultaneous splitter data Aug 11, 2015
10.05 Fo Splitter sim- higher int Tracks 3/12 - fiber data - simultaneous splitter data - int times increased to maintain light levels Aug 11, 2015
10.06 Fo Splitter sim - comp Tracks 5/10 - comparison of 3 days of Es/Ed head simultaneous data Aug 12, 2015
10.07 Fo Splitter sim- full track Tracks 5/10 - fiber data - simultaneous splitter data - USING FULL TRACK Aug 19, 2015
Day 10 and 11- Track 3/12- FO splitter test - Laser- MOBY Es/Ed head simultaneous w/ Rotation - 350, 375, 425, 450, 500 and 600 nm
11.01 Fo Splitter sim fiber data - simultaneous splitter data - with lasers Aug 14, 2015
11.02 saddle shapes Looking at changes in saddle shape with ES head rotation Aug 18, 2015
11.03 saddle shapes w/ lamp Looking at changes in saddle shape with ES head rotation - with lamp data added for comparison Aug 18, 2015
Day 12- Track 3/12- FO splitter test - Laser- MOBY Es/Ed head simultaneous w/o Rotation - lots of laser wavelengths
12.01 Fo Splitter sim - comp Fiber data - simultaneous splitter data - with lasers - comp with day 9-12 Aug 18, 2015
12.02 Fo Splitter sim - day 12 only Fiber data - simultaneous splitter data - with lasers Aug 18, 2015
12.03 Day 3 feature check Using laser data 715-720 nm and 920 nm and 955 nm to check if day 3 feature is from infrared Aug 18, 2015
12.04 400 nm int check 400 nm - integration time linearity check Aug 18, 2015
12.05 Fo Splitter sim - clean Fiber data - simultaneous splitter data - with lasers - comp with day 9-12 - CLEANED up verions 12.01 Aug 18, 2015
Day 13- Track 3/12- FO splitter- Laser and Oriel 60000 - MOBY Es/Ed head simultaneous - with Polarizer
13.01 Polarized Polarizer data - Es head Aug 19, 2015
13.02 Fo Splitter sim Simultaneous splitter data - with Oriel lamp and polarizing filter Aug 20, 2015
Day 14- Track 3/12- FO splitter - Oriel Lamp- MOBY Lu head simultaneous - Polarizer data - day14 = long FO
14.01 Polarized Polarizer data - Lu head Aug 20, 2015
14.02 Polarized - 10 nm Polarizer data - Lu head - 10 nm wavelengths Aug 20, 2015
14.03 Fo Splitter sim Simultaneous splitter data - with Oriel lamp and polarizing filter Aug 20, 2015
Day 15 - Track 3/12- FO splitter - Oriel Lamp- MOBY Lu head simultaneous - Polarizer data - day15 = short FO
15.01 Polarized Polarizer data - Lu head Aug 21, 2015
15.02 Polarized - 10 nm Polarizer data - Lu head - 10 nm wavelengths Aug 21, 2015
15.03 Polarized compared Comparing day 14 (long FO) and day 15 (short FO) data Aug 21, 2015
15.04 Effect on MOBY Polarization effect on MOBY data (M252-53) Sep 8, 2015
15.05 Mike Doc MOBY259 Pre-Deployment Calibrations - polarization tests Sep 28, 2015
Day 16 - Track 3/12- FO splitter - Oriel Lamp- MOBY Lu head simultaneous - 2 Polarizers crossed
16.01 crossed polarizers 2 Polarizers crossed - plots of data Aug 24, 2015
Day 17 - Track 3/12- FO splitter - Oriel Lamp- MOBY Lu head simultaneous - Polarizer data with Thorlabs Quartz Depolarizer behind Lu window
17.01 Depolarized With a depolarizer installed in the MOBY Lu heads window Aug 25, 2015
Day 18 - Track 3/12- FO splitter - Oriel Lamp- MOBY Lu head - Carols long-pass & short-pass filters filters on Track 12
18.01 Filter fun Looking at lamp and HgA lamps with filters Aug 28, 2015
Day 19-20 - Track 3/12- FO splitter - Laser - MOBY Lu head - Carols long-pass & short-pass filters on Track 12
19.01 Filter fun with lasers Looking at lasers with filters at may wavelengths Sept 1, 2015
19.02 2dark-5lite-2dark Looking at the data with more darks and lights averaged Sept 1, 2015
Day 21 - Track 3/12- FO splitter - Laser - MOBY Lu head - LP and SP on Track 12 - different combinations of Horizontal readout rates and Vertical shift speeds
20.01 Readout rates and shift speeds Looking at different combinations of Horizontal readout rates and Vertical shift speeds Sept 2, 2015
20.02 Sumary graphs (sort of) Some if the interesting things we have seen in the Horizontal readout rates and Vertical shift speeds data Sept 3, 2015
20.03 PDF Mike document on Andor camera shift speeds (rev 2) Apr 21, 20156
       
 
DATA PLOTS that cover many days
100.01
Darks/backgrounds
Darks over time
Aug 24, 2015
       

Camera details for BS01 and the software used during this experiment

BS01 camera, Andor iKon-M mod. DU934P-BR-DD, S/N: CCD-17878, I was running Andor SOLIS software: SOLIS ver. 4.23.30002.0,  SDK ver. 2.95.30002.0, installed via CD no. 2208071, copyright 2010, on a Windows 7 Pro/SP1 PC.

DAY 1 EMAIL from Mike
Hello Stephanie, you have data, @ /ftp1/Mike/HI-2015-03.

I ran the BS01cfg02 tonight, with the Ekspla(cfg09).
The BS01 returned from Resonon with a different Andor camera than it shipped out with -
I'll look for this new camera's spec/docs tomorrow.
Tonight was only a test to get some quick data for Ken's NOAA/JPSS talk.
I did not collect no stink'n Aux temp/humid data.
I'll try to setup the Aux data acquisition tomorrow,
and we'll see what more data Ken might like to have ASAP.

Tonight I scanned:
450 nm w/ all 14 FOs, then only FO #7, the only FOs #4,7,10,
350 nm w/ only FOs #4,7,10,
680 nm w/ only FOs #4,7,10.

Luck, MF
DAY2 EMAIL from Mike
Hi Steph,
There are some new data for you at the nas ftp. I changed the physical setup tonight, but only manipulated one parameter in software. I did not get to setup the Aux data acquisition.

There was only FO #7 coupled to the integrating sphere, without the FC ferrule coupler, and using a thin one-hole "jig" at the sphere's exit port - seen in photos 2015-07-01-01 to 05.jpg.

I de-coupled the Ekspla FO from the entrance port of the sphere - photos #06 to 08.  This was to keep the laser power up, and stable, but decrease the level at the sphere exit port, to keep the camera's exposure time highish.

I also monitored the sphere's power level via a PhotoDiode Monitor (PD Mon on the logsheet, in current units of nA) - photos 09,10,11 show the PD Monitor on the sphere, the Thermal Electric Controller (TEC), and the Keithley DVM (in "I" = current mode).

In software I only changed the "Shutter open & close delay" time (msec). I scanned background (b* files), then signal (s* files), then Stopped the laser to get a dark scan (d* files).  This repeated for shutter delays at 10,15,20,25,30,35 msec (open & close delays kept equal).

I did some rudimentary data-processing in MATLAB after each dark scan, - while the laser was re-started and stabilizing - and it looks to me like the shoulder at the high-pixel-side of the peak went away at longer shutter delay times.  I think the darks looked same as the backgrounds (at least near the laser peak).

Maybe tomorrow we'll try leaving shutter delay alone and change the integration time - so that the difference gets larger...
MF
DAY3 EMAIL from Mike
HI StePh,
Mark reminded me of a more interesting idea for data tonight: a sun scan ! There are photos of the setup, and of the sky conditions at your nas ftp, as well as a one-page log sheet.  There were 8 signals + 8 backgrounds acquired.  I did not get the Aux data running.  Track #7 FO was coupled to a MOBY Ed head (post-MOBY257 EdBot = head # Ed42, via a long ROV FO between the tent & the laser van), all other track/FOs were capped. The first 4 scans had Ed pointing straight up, sitting next to the tent. Then I thought it might be nice not to "see" the tent, so the next 4 had the Ed mounted on a tripod, pointed towards the (lower) sun.
Luck, MF
DAY4 and 5 EMAIL from Mike
Hi Steph,
There are some more BS01 data for you at the ftp, from day04 & 05. Day05 was essentially a repeat of 04, except that I cleaned out a bad FC connector on the splitter for 05.  There was epoxy at the bottom of the "WEAKER" splitter FO - I tried to take photos before I chipped it out. The epoxy did not allow the WEAK FO to seat on day04, so I think those scans are bogus - hence the day05 repeat.

I used the Oriel incandescent source with the 12" Spectralon sphere. I ran the TEC photodiode monitor and hand-logged it's current output. I bought some new HACA FC-to-bulkhead connectors and made a jig for the 12" sphere - see photos.  First I connected BS01 FO#7 to the jig/sphere - collected 5x sig/bac scans.  On the logsheet are photodiode monitor levels, ~19 uAmps, which we may need to normalize the BS01 data to.  Then I connected the single end of the splitter to the jig/sphere, and in-turn each of the split FO ends to BS01 FO#7 using the new FC-to-FC connector ADAFC1 - see photo of this conn.  I do not know what the loss off this connector is - I suppose we're about to find out !  One of the two split FO ends is labeled "WEAKER" the other is not labeled.  At the very end I collected 3x sets of sig/bac after the lamp was turned off * these signals are d* files, or darks.

I think from this set we can do a first check of the relative throughput of the two splitter ends.  I think they will be less than ~50%, as the splitter spec sheet says, because of losses at the FC-to-FC connector.

I thought the Oriel incandescent would be more stable.  Maybe the pd monitor levels will help correct for it's instability.  I can make a HACA jig for the NIST OL455 - it is 1/2 done - and we can repeat this with that source if needs be.  The OL455 also has a pd monitor, but it has no TEC / temperature control.

Luck !  MF
DAY6 EMAIL from Mike
Hi Steph,
I re-ran the day 05 splitter kine tonight. This is log sheet #1 of 2 for 29-Jul / page #6 overall. I only scanned the 2 bare splitter fibers - with ~50kADU @ splitter FO#2, then FO#1 at the same exposure & gain - to see if we can repeat the FO#2 / FO#1 ratio from day05.  There are again photodiode micro-amp readings to use for normalization.  These were sig & bac files, 5 each for the 2x FOs, for a total of 10 files, plus 1 pair of dark & bac @ the end.

The second setup was per Ken's suggestion. I got the Es head from MOBY257, and the EdBot fiber, and used the Oriel incandescent lamp to illuminate Es.  This is log sheet #2 of 2 / pg.7. The Es head & EdBot fibers' FO connectors were cleaned.  The Es-to-FO connection got coupling gel.  The Es head collector was also cleaned. I used the Es head's alignment jig for FEL cals to mate to the Oriel lamp. There are 6x photos of this setup.  The splitter input FO went into the end of the EdBot FO from the Es head.  Either splitter output FO#2 or FO#1 went into the BS01 FO#7.  I rotated the Es head 3 times for each splitter output FO connection, such that the Es head orientations were 1 = straight "up", 2 = 90 degree rotate, 3 = 180 deg rotate, 4 = 270 deg rotate. There were no photodiode monitor levels for normalization. This was a quick setup.  I imagine that the mating between Es & lamp was not perfect for each rotated state of the Es head.  I don't know how much the lamp varied during these measurements, so I don't expect the ratios to be flat & equal to one.  Each of the 4x rotations for the 2x FOs had a sig & bac file acquired, for a total of 16 .fits files.

Cheers, MF

DAY7 EMAIL from Mike
Hi Steph, guess what ?  Yup, there are more data at your ftp site !

There are 11 photos and 1 log sheet (pg.8) for day 07 = 30Jul2015 GMT, which was: MOBY257 Es head #Es40, via EdBot FO #3005, and splitter SN: CERL15050101, with MOBY Es & FO connected to splitter input FO, then 2x splitter output FOs connected to 2x BS01 input FOs: split FO#1 to BS01 FO#5, split FO#2 to BS01 FO#10. Measured were 10 rotation angles of the Es head @ 0:10:90 degrees - 0 deg was Es pointing directly at the light source, 90 deg was Es seeing the light source just grazing across the cosine bezel edge.
Each angle had 2x .FITS files collected, sig & bac, for a total of 20x files.

Tonight I did not try changing the bend in the splitter fibers - maybe tomorrow...

MF

DAY 8 EMAIL from Mike
Hi Steph,

And, I have another round of data from day08 - I repeated the 0-to-90 degree Es rotation of day07, butt with a longer MOBY FO cable between Es & the splitter. I was reading that the mode order is a function of fiber length, so it might be interesting to compare a short fiber = day7 via a ~3.75 m length MOBY EdBot FO versus day8 with a ~12.1 m length MOBY LuTop FO - both coupled to a MOBY Es head at one end, and coupled to the input of the splitter at the other end. I think the data processing for day8 would be ditto that of day7, ...including my above request for saddle overplots along the spectral dimension. There are 20x .FITS files and 1 log sheet .jpg at the ftp site for you!

Be thee here fore-warned - I have dug out our polarizer filters, and I have read up on the polarization tests we did for MOS, and I have read aboot Stan Hooker's polarization tests during SIRREX-7, and how polarization was characterized for POLDER, so I am planning
to collect some polarization data next week!

Thank you!  MF

DAY 9 EMAIL from Mike

Hello,

On Monday I got the BS01 auxiliary data cabling / acquisition running, and on Tuesday I re-ran the Day8 setup with these changes:
A.) increased exposure time,
B.) changed exposure time for different Es rotation angles to keep full ADU levels,
C.) changed fiber / track input at BS01 fiber bundle.

Attached here are my notes / figures until Steph gets back: <Day09_Results(MF).pdf>

I think the next setup will be using the laser source, and I heard Mike K's good idea
of fein-rotations at 0 to 12 deg.

Aloha, MF


Hi Steph,

I am in the process of FTPing to you the Day09 data for BS01 testing. Yesterday I sent a PDF with results of my attempt at data processing. There are 7x new photos: 2015-08-05*.jpg, there are 4x new data log sheets: *log_10...13.jpg, plus the writeup: Day09_Results(MF).pdf, and there will be 80x data files: s,b20150805*.fits.

One thing I messed up: I did not save a b (background) file for #10 i.e. b2015080510.fits, so I made one by copying from b2015050809.fits, which was at the same exposure time...

Once this is pau I'll be moving on to making day10 data ... after lunch !

Thank You, MF


I almost forgot...
I got the aux data stream for BS01 running again for day09. I just FTPed the file to the day09 folder = aux2015080501.txt I made it look like the files for HI-2015-01 ( see aux_descrip.txt ) but I added a 7th channel to log a thermistor at the splitter (Mark's request) (see day09 photo 2015-08-05-03=aux-splitter-therm.JPG).

This Thermistor is a Cantherm CWF-4B-103G3380, (which we used with the FISH for some kine external temperature), for which I converted Ohms to degrees in MATLAB via >> t5=mythermcnv(o5,'\wrk\thermist\thrm_Cantherm.dat'); for which you'll want the 3rd order polynomial coefficients in the attached thrm_Cantherm.dat.

FYI: I'm also attaching the R-vs-T doc for this thermistor:
Cantherm_RT.txt

Good Luck !  MF

DAY 10 EMAIL from Mike

Hi Steph,

Yesterday I started the experiment of a laser source input, to a rotating Es collector, to long MOBY fiber, to splitter, to BS01. I used split FO#1 to BS FO#3 / split FO#2 to BS FO#12, like day9 end. There are 3x log sheets, for 3x wavelengths = 350, 375, 400 nm,
plus 10x setup photos, with 10x sig&bac for each 0:10:90 deg rotation for each wavelength = a grand total of 60x FITS files. Oh yah, there is also an aux2015081301.txt file. On the log sheets are noted the internal laser PhotoDiode reading (Euv), and the reading for the TEC-PD on the sphere - I don't think you should worry about trying to normalize Andor data to these.

These first 3x near-UV wavelengths went very slowly because of the low laser power and hence long exposure times, but I expect to roll through more waves tonight - hopefully all remaining waves... Next on the list is a bare fiber, rotating through its' ± 12 deg acceptance angle, with laser input...

See YOu, MF

DAY 11 EMAIL from Mike

Hi Steph,
Day11 data are at your ftp site. 80 fits files + 1 aux file + 4x log sheets = page 17-to-20. These would be laser waves 425,450,500,600 nm. I found a BCJ email from 07-Aug-2015 suggesting "We only need a couple of waves past pixel 300 but if the scanning could be done more finely below pxl 300 we would map out the splitter ratio there." So maybe we'll want some more waves below pix 300...
MF

DAY 10/11 Thoughts EMAIL from Mike

Hi Steph,
R.E. Your fig.10 from pg.11.01 = Fo Splitter sim, (attached: plt_fo_splitter_mobyheads_day1011_t312__10.png)

1.) I'm beginning to think that rotation angle of Es collector is not an issue.
1.a) To me the saddle shape is an indicator of fiber mode behaviour (?)

2.) I'm beginning to think the track ratio is a function of track responses.
2.1) Q: Why does the track ratio not repeat for broadband vs monochrome source?
2.1.a) There were different exposure times / ADU levels for broad vs mono.
2.1.b) Different sources= ~point source / incandescent vs diffuse sphere / laser.
2.1.c) Stray light - SF thinks this is small except for 2nd order. etc.

I will repeat the 7x laser wavelengths tonight, to check repeatability of the track ratios, but only at a zero-degree Es rotation, probably filling in some additional wavelengths. Steph reminded me to add the Mike Kehoe IR artifact waves at "715-720 nm and 920 nm and 955 nm".

I will vary the exposure time, and/or the laser power - probably only at one wavelength - to see if the track ratio changes with ADU level (i.e. linearity).

R.E. the bare FO ± 12 degree rotation: I don't think this will work with an integrating sphere source, because the FO will be viewing different surfaces of the sphere as it rotates.  I think this requires a ~point source.  Thoughts?

Terrence is assembling the next MOBY and soon I will need to move to MOS/MOBY cals, so my time for the BS01 will be much reduced.
A.) I know the polarization check is on tap.  I think this should be somewhat straightforeward (ha).
B.) I have some bandpass filters/mirror that Mark bought from Thorlabs/Edmond, plus SMA & FC collimators from the FISH, but I have no clear plan r.e. how to go about testing their effects.

Aloha, MF

DAY 12 EMAIL from Mike

Hi Steph,
Day12 data are ftping now...

For Day12, I did not change the collector nor FO coupling from Day11. Es viewed the Ekspla integrating sphere.  Es was at zero deg rotation to said sphere, and its' rotation was not changed.  I re-ran the day10,11 laser wavelengths, and filled in some inbetween waves.  I reckon we should see if the day12 track ratios overplot with day10,11... I also did a short check at 400nm of ~linearity = varied only exposure time while keeping same source level to have a range of ADU signals to check for any affect on track ratio.  This is a cheap combination check - there could be a non-linearity in exposure time, and/or there could be a non-linearity in count level...  I also ran some near-IR waves for Mike Kehoe.  I saw no sign of signal at 30 sec exposure time.  This might be better re-run with a bare FO and/or a radiance head, because the Es diffuser in front of the sphere is already a somewhat insensitive sensor...

Luck, MF


DAY 13 EMAIL from Mike

Hi Steph,
Day13 data are heading your way. Only 1 log sheet scan = pg.23, plus 1 aux file = aux2015081901.txt.

The fits 2 pairs - sig&bac - of .fits files are Constant Wavelength, CW, lasers @ 543.5 & 632.8 nm.

The rest of the .fits files are via the polarizer, which is a filter on a rotation stage, illuminated by the Oriel incandescent source via the Spectralon sphere. They were all at the same 30 sec exposure time, so I only took a bac scan every 5th file.  I really needed to to scan at 60 sec or more but I was running late, and I'm sure this will be repeated...  The polarizer was rotated from zero to 90 deg at 10 deg intervals.  The photodiode monitor on the sphere had micro-Amps logged on the log sheet - hopefully we don't need to normalize to these...

I'm just now noticing that 1/2 of the .fits files are 4 MB size...

I took some photos.  The first 5 are of the CW lasers. The first 4 of the polarizer photos were to document how I aligned the filter in the rotator.

I have some docs I'll gather for the polarizer. I have the doc of the SIRREX7 polarizer test, which has a processing "parameter" which I think we should compute.  I'll get those together & send them next.

Thanks, MF

Here's the pages from the Melles catalog for our polarizer.
<attached= MellesGriot-DichroicSheetPolarizer.pdf>

Hi Steph,
Here are the pages from SIRREX7, chapt. 7.3, r.e. polarization
<attached= SIRREX7-Polarization.pdf>

I'd like to play with the "polarization parameter, P"
from section 7.3.3 and compare with their Table 11 results...
MF

DAY 14 EMAIL from Ken with Polarizing graph

Hi Steph,

Here is a graph with what one would expect to see (of course noise free). Notice, if you want, that the cosine is squared, due to Malus’s Law. Because of the trig identity cos(u)^2=(1+cos(2u))/2 you will also see this written as a function of cos(2*angle)….same thing. In fact when both MODIS and VIIRS polarization sensitivity were being measured, the clue that something was wrong was that they ended up with factors proportional to cos(4*angle) and not cos(2*angle). In general, you could take the raw data and do a fourier analysis of it, and the polarization sensitivity should show up as a cos(2* angle) component...

Ken

DAY 13-14 EMAIL from Mike with Polarizing filters and there use

Hi Ken,
The polarizer is a Melles Griot 03-fpg-007 - I attach here the "specs" from Melles <attached= MellesGriot-DichroicSheetPolarizer.pdf>

There are 2 "engravings on the outside edge" of the filter. I aligned the one near the Melles part# engraving with zero degrees on the rotator.  The rotator ended up aligned to the Lu head at 90 deg - i.e. zero deg on the rotator was toward & along the axis of the fiber connection to the Lu head. I took some photos of this setup (see Steph's "Photos" webpage, Date: 19-Aug-2015).  I'll have a look today at the min/max extinction of reflectance off Snug harbour water relative to the engravings / rotator / Lu head.

Rodger the depolarizer.  I'll get some help from Mark to place it in front of the Lu collector head's mirror...

MF

DAY 16 EMAIL from Mike

Hi Steph,
On Friday I ran the BS01+splitter+shortFO+Lu via sphere+incandescent through 2x polarizers.  The idea was to check the polarization ability of the optics to extinguish all signal when the optical axes were crossed. The Melles Griot specs say the polarizers are good for 350-650 nm range, so one might expect their ability to polarize light past 650 nm to decrease. I'd been wanting to make this measurement because Bob Barnes told the story that some NASA peeps had NOT done it, but I did not really understand why one should do it...

There was really only 2 net-signals measured, with the 2 optical axes parallel and with the axes crossed.  I had to increase the exposure time when they were crossed.  The interesting thing I saw was Mike Kehoe's IR artifact pop out with the polarizers crossed!

There is aboot to be ftp data for you = 1 log sheet, 6 photos, and 4 fits s/b pairs, plus one dark scan at the end with the lamp off. I'm curious how the dark compares to the previous background, because even though I try to baffle out all room light I have 2 laptops running and it is a small room...

On Monday I plan to open the Lu head and put the depolarizer behind the Lu window, then re-scan the single polarizer...

Thanks, MF

DAY 17 EMAIL from Mike

Hi Steph,
There are day17 depolarizer data coming your way. There are 10x sig/bac .fits pairs + one dark at the end. There is one log sheet scan = pg.27, + 2x Thorlabs*.pdf docs for the depolarizer. There are 8 photos.  I tried to document the orientation of the depolarizer - it has a flat edge on the filter to mark it's optical axis (see Thorlabs_DPU-25-A-AutoCADPDF.pdf) which I tried to orient at 45 deg to the plane of the beam to/from the mirror, for the flat face of the depolarizer facing the mirror (as per Ken's email suggestion).

MF

depol_aligndepolarizer_alignment

DAY 14-15 EMAIL from Ken on Polarization

HI Steph,

See if the attached document makes sense on how to use the fitted parameters (M12 and M13) for predicting the change in the MOBY measurement of Lu….then the correction would invert this….

ken

DAY 18 EMAIL from Mike

Hi Steph,
I took a few scans tonight = Day18. We are moving on to testing the long-pass & short-pass filters that Carol found-&-bought from Semrock. I think I forgot to
ftp this, so attached are some Semrock docs

BLP01-355R_Spectrum.txt, FF01-720_SP_Spectrum.txt
BLP01-355R-25_SemrockFilterDetails.pdf
FF01-720_SP-25_SemrockFilterDetails.pdf
Semrock_filter_documents_aug2015.pdf

and two plots I made from the .txt spectra

semrock_BLP01-355R.png, semrock_FF01-720_SP.png

The filters were mounted in Thorlabs 1 in dia tubes, then the two tubes were mounted together, then a Thorlabs collimator F220FC-A was mounted at the input & output ends. Docs for the collimators & the anti-reflective coating I did ftp to the doc dir. There are photos of the filters and their mounts.

The filters+collimators went between the splitter FO#1 and the BS01 track#16 FO. The splitter FO#2 was not filtered & was left connected to BS01 trk#3. The Lu head
and short FO connection was left unchanged from Day17, but I removed the depolarizer from the Lu head.

I thought I'd see if the filters removed the IR ghost which we saw via the crossed polarizers - they did - this was the first few three fits scans.

Then I tried to see if the filters removed any 2nd order signal from the UV lines from a HgA wavecal lamp. This idea came from the attached doc

Hg-Ar_notes(17Jul2014).txt

I think I forgot to ftp this too... This did not happen. Something in Lu head - maybe the mirror - and/or the fiber already removed the UV lines - just like I noted in the Hg-Ar...txt doc for MOBY... The HgA scans in fits files *.04 were saturated, but the *.05 were not. Maybe the 05 will be good for wave cal training!

Next / tomorrow I reckon will be laser line scans near the filters cut-on/off wavelengths...

Luck, MF


EMAIL from Carol

FEL lamp polarized output - document


DAY 19-20 EMAIL from Mike

Hi Steph,
There are some new BS data for you! I am combining day19/Friday with day20/Monday, because they are all Ekspla laser data through the combined LongPass + ShortPass Semrock filters, which are on Trk#12, with no filters on Trk#3 for reference.

The wavelengths scanned were a bit mixed up - sorry. On Friday there was no trk#12 sig below 360 nm. Monday I fein-scanned 357 & 362 nm = around the tiny 360 nm sig, then I scanned 410 nm to off-the-array 700 nm, on Monday.

I also tried to collect some data for your comparison of SL vs MOS. At 460 nm I got 2bac / 5sig / 2bac.  The file numbering might be a hassle for you - let me know if you think of a better way to number this kind of scan set...

There are 2x log sheets + 2x aux files + no new photos.

Good Luck!
MF

P.S. Terry is making progress assembling MOBY259
(I think we're waiting on Mark who's working on the new controller)
so this BS should be nearing a pause someday soon...


DAY 21 EMAIL from Mike

Hi Steph,

for the plot of SL vs MOS I collected some more 460 nm scans tonight, all at 2 sec exposure and 4x preamp gain, but at different combinations of Horizontal readout rates and Vertical shift speeds.  There are a bunch of dark scans at the end of it all.  There is one log sheet, one aux file.

Luck!  MF


DAY 21 EMAIL from Ken

HI Mike (and all),

Nice write up Mike.  I just read through the camera manual (finally) to see what it said.  A couple of things I think I can see from combining Mike’s writeup and Stephs page 20.2 (and reading of the manual and the Andor website).

1) we never want to use the 5MHz setting unless it is just for a quicklook.  In the specifications sheet for the camera (IKON-M) it says:
"5 MHz is for focusing/visualization mode only. “

2) 50kHz is too slow, as Mike points out this is slow enough, at least at the temperatures we are running the camera now, to have significant accumulation of dark counts on the array during readout. ( a factor of 20 slower than 1Mhz).

3) the funny fall off’s on edges occur for slow vertical transfer speeds and are worse for slow readout.  My first impression would be an uneven temperature distribution across the array…not as hot on the edges, so when you are taking along time to read the array out (slow readout speed and slow vertical transfer) you get this funny shape.  So we want to stay away from V = 64usecs.

4) Steph’s figure 7 seems to show that the horizontal readout at 3MHz has a spreading to higher wavelength pixels.  Now in the vertical direction, the website says to watch out for too fast a time, as the reduces the charge transfer efficiency and can cause some smearing.  We are doing vertical transfers limited to 4usec.  If you think about it , 3MHz is a horizontal transfer speed of 0.3 us.  It maybe that the camera really doesn’t do 3MHz very accurately, or not well enough for our application.   It looks like the 5MHz works better, so I maybe totally off….but this would point me to using the 1MHz rather than the 3Mhz..unless we figure out a different reason for that smearing.

5) I am partial to 1MHz, 4.25us vertical transfer….it has a low mean, it has a low std, it doesn’t show the smearing or asymmetry of the peak, it only takes 5 s for a  full image collection.  Probably faster as the 5 s number are with a 38us vertical shift speed rather than 4us.

Ken